OpenStudy (anonymous):

How do evaluate the following Intergral: (1,4)∫ e^√t/√t dt?

OpenStudy (anonymous):

try \(u=\sqrt{t}\)

OpenStudy (anonymous):

I got upto 2∫(1,4) ue^u/u du

OpenStudy (anonymous):

\[2\int_1^2e^udu\]

OpenStudy (anonymous):

via \[u=\sqrt{t}, du =\frac{1}{2\sqrt{t}}dt, 2du=\frac{1}{\sqrt{t}}dt, u(1)=1, u(4)=2\]

OpenStudy (anonymous):

how did u get \[\int\limits_{1}^{2}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

and what happens to \[\sqrt{t}\] that was at the bottom

OpenStudy (anonymous):

nm i get it ... thank u